# Few-Shot Adversarial Low-Rank Fine-Tuning of Vision-Language Models

### v2

@@ Abstract
Vision-Language Models (VLMs) such as CLIP have shown remarkable performance in cross-modal tasks through large-scale contrastive pre-training. To adapt these large transformer-based models efficiently for downstream tasks, Parameter-Efficient Fine-Tuning (PEFT) techniques like LoRA have emerged as scalable alternatives to full fine-tuning, especially in few-shot scenarios. However, like traditional deep neural networks, VLMs are highly vulnerable to adversarial attacks, where imperceptible perturbations can significantly degrade model performance. Adversarial training remains the most effective strategy for improving model robustness in PEFT. In this work, we propose AdvCLIP-LoRA, the first algorithm designed to enhance the adversarial robustness of CLIP models fine-tuned with LoRA in few-shot settings. Our method formulates adversarial fine-tuning as a minimax optimization problem and provides theoretical guarantees for convergence under smoothness and nonconvex-strong-concavity assumptions. Empirical results across eight datasets using ViT-B/16 and ViT-B/32 models show that AdvCLIP-LoRA significantly improves robustness against common adversarial attacks (e.g., FGSM, PGD), without sacrificing much clean accuracy. These findings highlight AdvCLIP-LoRA as a practical and theoretically grounded approach for robust adaptation of VLMs in resource-constrained settings. The code is available at https://github.com/sajjad-ucsb/AdvCLIP-LoRA.
@@ 1 Introduction
Vision-Language Models (VLMs), such as CLIP [1], have become foundational in learning cross-modal representations by aligning visual and textual embeddings through large-scale contrastive pre-training [2, 3, 4]. While these models enable effective zero-shot and few-shot adaptation [5, 6], their larger transformer-based variants [7] demonstrate superior performance (e.g., CLIP’s ViT-L/14 surpasses ViT-B/16 by over 6% on ImageNet [8]). However, these large models typically contain billions of trainable parameters, making full fine-tuning (FFT) computationally expensive and inefficient, particularly for task-specific adaptations. To address this, Parameter-Efficient Fine-Tuning (PEFT) methods have gained traction, particularly in NLP, where techniques like adapters [9, 10, 11] and prompt tuning [12, 13] reduce overhead, by adding a small number of trainable parameters or trainable prompt tokens while keeping the rest of the model frozen.
@@
Among PEFT methods, Low-Rank Adaptation (LoRA) [14] offers an efficient alternative by fine-tuning only low-rank matrices, enabling single-GPU adaptation of billion-parameter models [15] while matching full fine-tuning performance [14]. Recent work by [16] employed LoRA in the context of few-shot VLMs, demonstrating improved accuracy across various tasks and models. Unlike few-shot prompt tuning [17, 18, 6], which involves computationally intensive optimization of textual prompts, or adapter-based methods [19, 5] that often demand extensive hyperparameter tuning [20], LoRA provides a more scalable and portable solution for fine-tuning VLMs [16].
@@
Despite their impressive capabilities, VLMs share the susceptibility of traditional deep neural networks (DNNs) to adversarial attacks, where imperceptible perturbations can significantly degrade model performance [21, 22]. This vulnerability is particularly concerning in the visual domain, where adversarial noise can be more subtle and difficult to detect compared to textual modifications. Extensive research in computer vision has demonstrated that adversarial training remains the most effective approach for developing robust DNNs resistant to adversarial perturbations [23]. When applied to PEFT paradigms, this adversarial training is typically implemented during the fine-tuning phase rather than during initial pre-training. More recently, studies [24, 25, 26] have explored few-shot prompt tuning as a means of adversarial adaptation. For instance, [25] trains the clean text embedding with the adversarial image embedding to improve adversarial robustness. However, despite LoRA’s established effectiveness for standard fine-tuning tasks, its potential for enhancing adversarial robustness in VLMs remains largely unexplored. This work addresses this gap by introducing AdvCLIP-LoRA, a novel algorithm designed to enhance the adversarial robustness of CLIP models fine-tuned with LoRA in few-shot settings.
Before delving into the details, we summarize our main contributions. 1) We propose AdvCLIP-LoRA, which, to the best of our knowledge, is the first algorithm designed to enhance the adversarial robustness of CLIP models fine-tuned with LoRA in few-shot settings by formulating and solving a minimax optimization problem. 2) We provide theoretical guarantees by proving that, under assumptions of non-convex–strong-concavity and smoothness of the objective functions in our minimax formulation, the primal function defined as $\Phi(\cdot)=\max_{\delta\in\Delta}f(\cdot,\delta)$ converges to a stationary solution. 3) We conduct extensive experiments across eight datasets using ViT-B/16 and ViT-B/32 CLIP models to evaluate the effectiveness of our approach. Results demonstrate that AdvCLIP-LoRA significantly improves the robustness of LoRA-fine-tuned CLIP models against FGSM and PGD adversarial attacks in few-shot settings.
@@ 2 Preliminaries and Related Work
2.1 Alternative Strategies for PEFT
Prompt tuning has emerged as an alternative to weight tuning for parameter-efficient adaptation of VLMs. CoOp [27] learns continuous prompt tokens appended to class names, while CoCoOp [28] extends this by generating instance-specific prompts conditioned on images. Other variants, such as ProGrad [6] and KgCoOp [29], project prompts toward handcrafted templates to preserve pretrained knowledge. PLOT [18] jointly adapts both image and text modalities using an optimal transport objective, and MaPLe [30] further extends this by coupling interdependent prompts across vision and text encoders.
While prompt tuning is highly parameter-efficient, it can underperform in generalization [31] and scale poorly with larger datasets [16]. Moreover, as shown in [32, 16], prompt-based methods typically require significantly longer training times compared to weight-based approaches. Given that prompt tuning is conceptually orthogonal to our LoRA-based adaptation method, we do not include direct comparisons in our main experiments. However, as shown in Fig. 1(a), our method, AdvCLIP-LoRA, achieves comparable and in some cases superior clean accuracy to these non-robust prompt-tuning approaches and to standard CLIP-LoRA, despite our focus on adversarial robustness.
Adversarial Prompt Tuning. Several recent works attempt to improve the robustness of prompt-tuned VLMs. APT [24] learns robust text prompts via adversarial training, while FAP [22] leverages multimodal prompts and proposes a loss function that balances the connection between natural and adversarial features across modalities. However, these approaches suffer from reduced clean accuracy—sometimes performing worse than even zero-shot CLIP—whereas AdvCLIP-LoRA consistently outperforms them, as illustrated in Fig. 1(b).
(a) AdvCLIP-LoRA vs. Non-Robust PEFT Methods (CLIP-Adapter [19], PLOT++ [18], KgCoOp [29], TaskRes [33], MaPLe [30], ProGrad [6], CLIP-LoRA [16]).
(b) AdvCLIP-LoRA vs. Robust PEFT Methods (APT [24], AdvMaPLe [30], AdvVLP [34], FAP [34], ER-APT [26]).
Figure 1: $16$-shot comparison of AdvCLIP-LoRA with both non-robust and robust PEFT methods using the ViT-B/32 model across eight datasets. Results for non-robust and robust PEFT baselines are taken from [16] and [26], respectively.
@@ 2.2 Few-Shot Fine-Tuning for VLMs
In vision-language classification tasks, predictions are made by leveraging the pretrained alignment between visual and textual modalities. Given a label set of $K$ classes, one first constructs natural language descriptions, or prompts [35], denoted as $\{c_{k}\}_{k=1}^{K}$, where each $c_{k}$ is a textual phrase such as “a photo of a [class name].” These prompts are embedded using a frozen text encoder $\theta_{t}$, yielding normalized representations $\mathbf{z}_{k}^{(T)}=\theta_{t}(c_{k})\in\mathbb{R}^{d}$. Similarly, an image $\mathbf{x}_{i}$ is embedded via a visual encoder $\theta_{v}$ to obtain $\mathbf{z}_{i}^{(I)}=\theta_{v}(\mathbf{x}_{i})\in\mathbb{R}^{d}$, also normalized to unit length. The prediction logits are computed as the cosine similarity between each image-text pair. These logits are converted into a probability distribution over classes using a softmax with temperature scaling:
@@
where $\gamma$ is a softmax-temperature parameter. The predicted label for image $\mathbf{x}_{i}$ is the one with the highest posterior probability: $\hat{k}=\arg\max_{k}\,p_{i,k}.$ This form of zero-shot prediction directly mirrors the contrastive training setup used in large-scale VLM pretraining, such as CLIP [1], and allows models to generalize to novel classification tasks without any fine-tuning on the target domain.
To further adapt vision-language models to downstream tasks, the few-shot setting assumes access to a limited number of labeled examples per target class—typically fewer than $16$. Given $N$ such labeled support images per class, we denote the one-hot encoded ground-truth label for image $\mathbf{x}_{i}$ as $y_{ik}$, where $y_{ik}=1$ if $\mathbf{x}_{i}$ belongs to class $k$, and 0 otherwise. Classification probabilities $p_{i,k}$ are obtained as in the zero-shot setup, and the model is adapted by minimizing the cross-entropy loss: $\mathcal{L}_{\text{CE}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}y_{ik}\ln p_{i,k}.$
This adaptation can be implemented in several ways. One strategy is to optimize the input prompts $\{c_{k}\}_{k=1}^{K}$ directly—an approach inspired by prompt tuning techniques [18]. Alternatively, one may choose to fine-tune lightweight, task-specific modules such as adapter layers [19] or low-rank parameterizations like LoRA [16], leaving the backbone encoders frozen.
@@ 2.3 Fine-Tuning VLMs via LoRA
Low-Rank Adaptation (LoRA) [14] is a highly promising PEFT method, enabling efficient fine-tuning of large models by freezing the entire pre-trained model and introducing low-rank, trainable matrices within each layer. In LoRA, given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the weight update is achieved through a low-rank decomposition $W_{0}+\Delta W=W_{0}+BA$, where the training occurs on matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$, with $r\ll\min(d,k)$. The values in $A$ are initialized randomly via a Gaussian distribution, while $B$ is initialized as a zero matrix. This setup ensures that no low-rank update occurs before training, meaning that the output remains unchanged initially.
Although the original LoRA paper applies the low-rank matrices to the attention matrices of transformer-based architectures, [16] extends LoRA to all matrices in the vision and text encoders of VLMs. This adaptation leads to improved accuracy over prompt-based methods across various CLIP architectures and datasets [16].
@@ 2.4 Adversarial Robustness
$\hat{x}=x+\underset{\delta\in\Delta}{\arg\max}\;\mathcal{L}(h(x+\delta),y),$ (2)
where $\mathcal{L}$ is the training loss function. This formulation represents an optimization problem where the perturbation $\delta$ is chosen such that the classifier’s output is maximally disrupted while staying within a bounded set. Methods like Projected Gradient Descent (PGD) [23] are commonly employed to solve this optimization problem. Given the vulnerability of deep learning models to these perturbations [21], it becomes crucial to defend against such adversarial attacks.
One of the most effective strategies for defending against adversarial attacks is adversarial training, as proposed by [23]. When $h_{W}$ denotes a classifier parameterized by $W$, adversarial training seeks to solve the following minimax optimization problem:
@@
where $\mathcal{D}$ represents the underlying data distribution. This approach effectively trains the classifier to be robust against adversarial perturbations by simultaneously minimizing the classifier’s loss and maximizing the perturbation within a bounded set.
@@ 3 Proposed Algorithm
1:Input: Learning rates $\eta_{w}$ and $\eta_{\delta}$, batch-size $M$, number of iterations $T$.
2:Initialize: $A_{0}\sim\mathcal{N}(0,\sigma^{2})$, $B_{0}=0$.
3:for iteration $t\leftarrow 1$ to $T$ do
4:  Draw a collection of i.i.d. data samples $\left\{\xi_{i}\right\}_{i=1}^{M}$
5:  $\delta_{t}=\mathcal{P}_{\Delta}\left(\delta_{t-1}+\eta_{\delta}(\frac{1}{M}\sum_{i=1}^{M}\nabla_{\delta}F(W_{t-1},\delta_{t-1};\xi_{i}))\right)$ $\triangleright$ Update and project the perturbation.
6:  $A_{t}=A_{t-1}-\eta_{w}\left(\frac{1}{M}\sum_{i=1}^{M}\nabla_{A}F(W_{t-1},\delta_{t};\xi_{i})\right)$
7:  $B_{t}=B_{t-1}-\eta_{w}\left(\frac{1}{M}\sum_{i=1}^{M}\nabla_{B}F(W_{t-1},\delta_{t};\xi_{i})\right)$ $\triangleright$ Update the low-rank matrices $A$ and $B$.
@@
9:Randomly draw $\hat{A},\hat{B}$ from $\{A_{t},B_{t}\}_{t=1}^{T}$ at uniform.
10:Return: $\hat{A},\hat{B}$.
Figure 2: Refer to caption: Trainable Parameters, Refer to caption: Frozen Parameters. Illustration of AdvCLIP-LoRA algorithm. During each iteration $t$, the perturbation $\delta_{t}$ is updated and applied to the input image batch. Subsequently, the low-rank matrices $A$ and $B$ are optimized, while the rest of the model remains frozen.

### v3

@@ Abstract
Vision-Language Models (VLMs) such as CLIP have shown remarkable performance in cross-modal tasks through large-scale contrastive pre-training. To adapt these large transformer-based models efficiently for downstream tasks, Parameter-Efficient Fine-Tuning (PEFT) techniques like (Low-Rank Adaptation) LoRA have emerged as scalable alternatives to full fine-tuning, especially in few-shot scenarios. However, like traditional deep neural networks, VLMs are highly vulnerable to adversarial attacks, where imperceptible perturbations can significantly degrade model performance. Adversarial training remains the most effective strategy for improving model robustness in PEFT. In this work, we propose AdvCLIP-LoRA, to our knowledge the first method designed to enhance the adversarial robustness of CLIP models fine-tuned with LoRA in few-shot settings. Our method formulates training as a minimax optimization over low-rank adapters and adversarial perturbations, enabling robust adaptation with a small trainable footprint. Across eight datasets and two backbones (ViT-B/16 and ViT-B/32), AdvCLIP-LoRA achieves state-of-the-art performance in few-shot classification, adversarial base-to-new generalization, and cross-dataset transfer, delivering higher adversarial robustness than prompt tuning baselines without sacrificing much clean accuracy. These findings highlight AdvCLIP-LoRA as a practical approach for robust adaptation of VLMs in resource-constrained settings.
@@ 1 Introduction
Vision-Language Models (VLMs), such as CLIP Radford et al. (2021), have become foundational in learning cross-modal representations by aligning visual and textual embeddings through large-scale contrastive pre-training Jia et al. (2021); Li et al. (2022b); Yao et al. . While these models enable effective zero-shot and few-shot adaptation Zhang et al. (2022); Zhu et al. (2023), their larger transformer-based variants Vaswani (2017) demonstrate superior performance (e.g., CLIP’s ViT-L/14 surpasses ViT-B/16 by over 6% on ImageNet Deng et al. (2009)). However, these large models typically contain billions of trainable parameters, making full fine-tuning (FFT) computationally expensive and inefficient, particularly for task-specific adaptations.
@@
To address this, Parameter-Efficient Fine-Tuning (PEFT) methods have gained traction, particularly in NLP, where techniques like adapters Chen et al. (2022); Karimi Mahabadi et al. (2021); Rebuffi et al. (2017) and prompt tuning Jia et al. (2022); Li & Liang (2021) reduce overhead, by adding a small number of trainable parameters or trainable prompt tokens while keeping the rest of the model frozen. Among PEFT methods, Low-Rank Adaptation (LoRA) Hu et al. (2021) offers an efficient alternative by fine-tuning only low-rank matrices, enabling single-GPU adaptation of billion-parameter models Dettmers et al. (2023) while matching full fine-tuning performance Hu et al. (2021). Recent work by Zanella & Ben Ayed (2024) employed LoRA in the context of few-shot VLMs, demonstrating improved accuracy across various tasks and models. Unlike few-shot prompt tuning Bulat & Tzimiropoulos (2023); Chen et al. ; Zhu et al. (2023), which involves computationally intensive optimization of textual prompts, or adapter-based methods Gao et al. (2024); Zhang et al. (2022) that often demand extensive hyperparameter tuning Silva-Rodriguez et al. (2024), LoRA provides a more scalable and portable solution for fine-tuning VLMs Zanella & Ben Ayed (2024).
@@
Despite their impressive capabilities, VLMs share the susceptibility of traditional deep neural networks (DNNs) to adversarial attacks, where imperceptible perturbations can significantly degrade model performance Szegedy et al. (2013); Zhou et al. (2023). This vulnerability is particularly concerning in the visual domain, where adversarial noise can be more subtle and difficult to detect compared to textual modifications. Extensive research in computer vision has demonstrated that adversarial training remains the most effective approach for developing robust DNNs resistant to adversarial perturbations Madry et al. (2018). When applied to PEFT paradigms, this adversarial training is typically implemented during the fine-tuning phase rather than during initial pre-training. More recently, studies Li et al. (2024); Zhang et al. (2024); Jia et al. (2025) have explored few-shot prompt tuning as a means of adversarial adaptation. For instance, Zhang et al. (2024) trains the clean text embedding with the adversarial image embedding to improve adversarial robustness. APT Li et al. (2024) learns robust text prompts via adversarial training, while FAP Zhou et al. (2023) leverages multimodal prompts and proposes a loss function that balances the connection between natural and adversarial features across modalities.
Figure 1: Few-shot performance across datasets under clean and adversarial evaluation. Spider plots show top-1 accuracy (%) for Clean (top row) and PGD-100 (bottom row) on eight datasets at shot counts $\{1,2,4,8,16\}$ with ViT-B/32. Each polygon denotes a method (larger area is better).
Despite their effectiveness, adversarial prompt-based methods exhibit two limitations: (i) they often attain robustness by sacrificing substantial clean accuracy, especially in the extreme few-shot regime (1–4 shots), where many underperform even zero-shot CLIP (Fig. 1, top); and (ii) their robustness typically improves only as the shot count increases, with some methods struggling to gain robustness in the extreme few-shot regime (Fig. 1, bottom). Although LoRA has proven effective for standard fine-tuning, its use for enhancing adversarial robustness in few-shot VLMs remains largely unexplored. We address this gap with AdvCLIP-LoRA, which fine-tunes CLIP using LoRA adapters under a minimax objective. As shown in Fig. 1, our simple AdvCLIP-LoRA avoids the above trade-offs, delivering superior robustness and higher clean accuracy, consistently outperforming adversarial prompt-tuning baselines on both clean and PGD metrics for the majority of shots.
Before delving into the details, we summarize our main contributions.
  • We investigate LoRA for adversarially robust few-shot VLMs, a setting largely dominated by prompt-based strategies, and introduce AdvCLIP-LoRA, which frames adaptation as a minimax optimization problem and solves it efficiently.
  • We conduct extensive experiments on eight datasets with ViT-B/16 and ViT-B/32 backbones, covering few-shot classification, adversarial base-to-new generalization, and cross-dataset transfer; AdvCLIP-LoRA significantly improves robustness to strong attacks (e.g., PGD) in most settings with minimal loss in clean accuracy.
  • We present comprehensive ablations that analyze design choices and hyperparameters, providing guidance for practical deployment.
  • Under standard assumptions from the minimax optimization literature (e.g., smooth objectives and bounded gradients), we establish convergence guarantees for the primal function $\Phi(\cdot)=\max_{\delta\in\Delta}f(\cdot,\delta)$ to a stationary point, with rates matching classical results.
@@ 2.1 Few-Shot Fine-Tuning for VLMs
In vision-language classification tasks, predictions are made by leveraging the pretrained alignment between visual and textual modalities. Given a label set of $K$ classes, one first constructs natural language descriptions, or prompts Liu et al. (2023a), denoted as $\{c_{k}\}_{k=1}^{K}$, where each $c_{k}$ is a textual phrase such as “a photo of a [class name].” These prompts are embedded using a frozen text encoder $\theta_{t}$, yielding normalized representations $\mathbf{z}_{k}^{(T)}=\theta_{t}(c_{k})\in\mathbb{R}^{d}$. Similarly, an image $\mathbf{x}_{i}$ is embedded via a visual encoder $\theta_{v}$ to obtain $\mathbf{z}_{i}^{(I)}=\theta_{v}(\mathbf{x}_{i})\in\mathbb{R}^{d}$, also normalized to unit length. The prediction logits are computed as the cosine similarity between each image-text pair. These logits are converted into a probability distribution over classes using a softmax with temperature scaling:
@@
where $\gamma$ is a softmax-temperature parameter. The predicted label for image $\mathbf{x}_{i}$ is the one with the highest posterior probability: $\hat{k}=\arg\max_{k}\,p_{i,k}.$ This form of zero-shot prediction directly mirrors the contrastive training setup used in large-scale VLM pretraining, such as CLIP Radford et al. (2021), and allows models to generalize to novel classification tasks without fine-tuning on the target domain.
To further adapt vision-language models to downstream tasks, the few-shot setting assumes access to a limited number of labeled examples per target class—typically fewer than $16$. Given $N$ such labeled support images per class, we denote the one-hot encoded ground-truth label for image $\mathbf{x}_{i}$ as $y_{ik}$, where $y_{ik}=1$ if $\mathbf{x}_{i}$ belongs to class $k$, and 0 otherwise. Classification probabilities $p_{i,k}$ are obtained as in the zero-shot setup, and the model is adapted by minimizing the cross-entropy loss:
$\mathcal{L}_{\text{CE}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}y_{ik}\ln p_{i,k}.$ (2)
This adaptation can be implemented in several ways. One strategy is to optimize the input prompts $\{c_{k}\}_{k=1}^{K}$ directly, an approach inspired by prompt tuning techniques Chen et al. . Alternatively, one may fine-tune lightweight, task-specific modules such as adapter layers Gao et al. (2024) or low-rank parameterizations like LoRA Zanella & Ben Ayed (2024), leaving the backbone encoders frozen.
@@ 2.2 Fine-Tuning VLMs via LoRA
Low-Rank Adaptation (LoRA) Hu et al. (2021) is a highly promising PEFT method, enabling efficient fine-tuning of large models by freezing the entire pre-trained model and introducing low-rank, trainable matrices within each layer. In LoRA, given a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, the weight update is achieved through a low-rank decomposition $W_{0}+\Delta W=W_{0}+BA$, where the training occurs on matrices $A\in\mathbb{R}^{r\times k}$ and $B\in\mathbb{R}^{d\times r}$, with $r\ll\min(d,k)$. The values in $A$ are initialized via a Gaussian distribution, while $B$ is initialized as a zero matrix. This setup ensures that no low-rank update occurs before training, meaning that the output remains unchanged initially.
Although the original LoRA paper applies the low-rank matrices to the attention matrices of transformer-based architectures, Zanella & Ben Ayed (2024) extends LoRA to all matrices in the vision and text encoders of VLMs. This adaptation leads to improved accuracy over prompt-based methods across various CLIP architectures and datasets Zanella & Ben Ayed (2024).
@@ 2.3 Adversarial Robustness
$\hat{x}=x+\arg\max_{\delta\in\Delta}\mathcal{L}(h(x+\delta),y),$ (3)
where $\mathcal{L}$ is the training loss function. This formulation represents an optimization problem where the perturbation $\delta$ is chosen such that the classifier’s output is maximally disrupted while staying within a bounded set. Methods like Projected Gradient Descent (PGD) Madry et al. (2018) are commonly employed to solve this optimization problem. Given the vulnerability of deep learning models to these perturbations Szegedy et al. (2013), it becomes crucial to defend against such adversarial attacks.
One of the most effective strategies for defending against adversarial attacks is adversarial training, as proposed by Madry et al. (2018). When $h_{W}$ denotes a classifier parameterized by $W$, adversarial training seeks to solve the following minimax optimization problem:
@@
where $\mathcal{D}$ represents the underlying data distribution. This approach effectively trains the classifier to be robust against adversarial perturbations by simultaneously minimizing the classifier’s loss and maximizing perturbation within a bounded set.
@@ 3 Proposed Algorithm
Figure 2: Refer to caption: Trainable Parameters, Refer to caption: Frozen Parameters. Illustration of AdvCLIP-LoRA algorithm. During iteration $t$, the perturbation $\delta_{t}$ is updated and applied to the input image batch. Subsequently, the low-rank matrices $A$ and $B$ are optimized, while the rest of the model remains frozen.
@@
1:Training samples $\mathcal{X}$, batch-size $M$, learning rates $\eta_{w}$, $\eta_{\delta}$
[…]
4:for epoch $1\ldots T$ do
5:  for minibatch $M\subset\mathcal{X}$ do
6:   for $j=1\ldots\tau$ do
7:     $\delta_{t}=\mathcal{P}_{\Delta}\left(\delta_{t-1}+\eta_{\delta}(\frac{1}{M}\sum_{i=1}^{M}\nabla_{\delta}F(W_{t-1},\delta_{t-1};\xi_{i}))\right)$
@@
9:   $A_{t}=A_{t-1}-\eta_{w}\left(\frac{1}{M}\sum_{i=1}^{M}\nabla_{A}F(W_{t-1},\delta_{t};\xi_{i})\right)$ $\triangleright$ Update the low-rank matrix AA
10:   $B_{t}=B_{t-1}-\eta_{w}\left(\frac{1}{M}\sum_{i=1}^{M}\nabla_{B}F(W_{t-1},\delta_{t};\xi_{i})\right)$ $\triangleright$ Update the low-rank matrix $B$
11:  end for
12:end for